OpenStudy (anonymous):

what is the slope of the line:y-3=-3(x+4)

OpenStudy (smurfy14):

-3x+4=y-3 y=-3x+7 slope is -3

OpenStudy (anonymous):

thank u SO much smurfy14!:)

OpenStudy (anonymous):

oh wait!what is the slope of the line:y-4=3(x-6)

OpenStudy (smurfy14):

lemme see

OpenStudy (smurfy14):

3x-18=y-4 y=3x-14 slope is 3

OpenStudy (anonymous):

thx!u r a lifesaver!:)

OpenStudy (smurfy14):

lol :)

OpenStudy (anonymous):

ooh.i have another prob. multiply: (x-5)(x-4)

OpenStudy (anonymous):

please can u help?;)

OpenStudy (anonymous):

I can help you with that problem if you want

OpenStudy (anonymous):

ok.:)

OpenStudy (anonymous):

We basically just multiply the terms. By doing this we get: x^2-4x-5x+20, then we add like terms and get: x^2-9x+20

OpenStudy (anonymous):

any other cute problems?

OpenStudy (anonymous):

thx!!!!;)

OpenStudy (anonymous):

yah. multiply: 5c^3(3c^4-6c^2 -4c-5)

OpenStudy (anonymous):

never mind.

OpenStudy (anonymous):

oh, come on i had answer all typed up

OpenStudy (anonymous):

ok what is it?

OpenStudy (anonymous):

i thought u had the answer?:)?

OpenStudy (anonymous):

The answer is; 15c^7-30c^5-20c^4-25c^3

OpenStudy (anonymous):

Don't worry, i have the answers

OpenStudy (anonymous):

ok;)

OpenStudy (anonymous):

What time is it?

OpenStudy (anonymous):

9:51 pm

OpenStudy (anonymous):

you mean 11:52

OpenStudy (anonymous):

no.its 9:52 now

OpenStudy (anonymous):

your 2 hours behind

OpenStudy (anonymous):

multiply: (9x^2 +3)(9x^2 -3)

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

hi chloe, do you understand the slope equation of the line formula? y=mx +b the m here is your slope you are asking,,and the b=y intercept .....so any time you need to find a slope you need to do computation to make the equation as y=mx +b

OpenStudy (anonymous):

and heres another one. multiply:(4x+9)(5x-3)

OpenStudy (anonymous):

and another.multiply:(5x+8y)(5x-8y)

OpenStudy (anonymous):

you can solve (4x+9)(5x-3) by FOIL method, do you wanna try?

OpenStudy (anonymous):

and another.multiply:3x^3(9x^2+2y)

OpenStudy (anonymous):

and another.multiply:-5x^2y(x^2 -3xy+7y^3

OpenStudy (anonymous):

and another.multiply:-6x^2y(x^2+4xy+2y^3)

OpenStudy (anonymous):

im pretty sure that you can solve any math prob if you will try to do it first then ask questions if you get stuck...lol....

OpenStudy (anonymous):

:(?

OpenStudy (anonymous):

and lastly, multiply:(3x+y)(3x-2y)

OpenStudy (anonymous):

can someone please help me?